OpenStudy (anonymous):

What is the difference in minutes between the 2 routes when driving 25 mph. the bridge is 18 miles and the damn is 12 miles

OpenStudy (anonymous):

...incomplete...

OpenStudy (anonymous):

d=V*t t= d/V

OpenStudy (anonymous):

times will be in hours; find the difference and convert to minutes..

OpenStudy (anonymous):

so 42.6?

OpenStudy (anonymous):

whoops.

OpenStudy (anonymous):

t1 = 18/25 and t2 = 12/25

OpenStudy (anonymous):

sorry about that.

OpenStudy (anonymous):

brain seizure typo

OpenStudy (anonymous):

use units to determine these questions\[25\frac{miles}{hour}\] if you want time(hours) to end up on top then you have to divide by the rate of speed

OpenStudy (anonymous):

 If yuo want to solve for one, cover it up and the triangle helps you determine which operations to perform

OpenStudy (anonymous):

ok t1 18/25=n.72 t2=12/25=.48 difference =.24 (hr) .24x 60=14.4 C0rrect?

OpenStudy (anonymous):

yep.

OpenStudy (anonymous):

minutes.

OpenStudy (anonymous):

ty :)